# Data Augmentation for Continual RL via Adversarial Gradient Episodic Memory

### v1

@@ Abstract
Data efficiency of learning, which plays a key role in the Reinforcement Learning (RL) training process, becomes even more important in continual RL with sequential environments. In continual RL, the learner interacts with non-stationary, sequential tasks and is required to learn new tasks without forgetting previous knowledge. However, there is little work on implementing data augmentation for continual RL. In this paper, we investigate the efficacy of data augmentation for continual RL. Specifically, we provide benchmarking data augmentations for continual RL, by (1) summarising existing data augmentation methods and (2) including a new augmentation method for continual RL: Adversarial Augmentation with Gradient Episodic Memory (Adv-GEM). Extensive experiments show that data augmentations, such as random amplitude scaling, state-switch, mixup, adversarial augmentation, and Adv-GEM, can improve existing continual RL algorithms in terms of their average performance, catastrophic forgetting, and forward transfer, on robot control tasks. All data augmentation methods are implemented as plug-in modules for trivial integration into continual RL methods.
@@ 1 Introduction
  • We introduce a novel data augmentation method, Adv-GEM, specifically designed for continual RL. Our experimental results indicate that Adv-GEM achieves the highest positive impact on average performance compared to other augmentation techniques, setting a new state-of-the-art for continual RL in robotic control tasks.
@@ 2.1 Continual RL
The standard RL of a single task can be described as a Markov Decision Process (MDP) consisting of a tuple $\langle S,A,P,r,\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $P$ is the transition probability, $R:S\times A\to\mathbb{R}$ is the reward function to map the current state-action pair to a reward value, and $\gamma$ is the discount factor for mitigating the impact of long-term benefits. The final objective of RL agent is to maximize the discounted cumulative reward.
@@ 3 Data Augmentation for Continual RL
In this section, we introduce how to use data augmentation for continual RL. We start with the problem formulation in Section 3.1. Then, in Section 3.2, we introduce 8 different data augmentation methods that can increase the diversity of the training data for continual RL. Finally, we present the general data augmentation framework for continual RL in Section 3.3.
@@ 3.2 Data Augmentation Methods
Figure 1: Process for Adv-GEM data generation. For each previous task, we store samples in the episodic memory $\mathcal{M}$. While we optimize the current task, we sample data from the replay buffer $\mathcal{D}$. We generate the adversarial examples by embedding the gradient episodic memory, where $H$ is the Heaviside step function, $g_{\mathcal{D}}$ is the gradient of the policy network at the current task, $g_{\mathcal{D}}=\nabla_{s_{i}\sim\mathcal{D}}\mathcal{L}_{\pi_{\theta}}(s_{i})$, and $g_{\mathcal{M}}$ is the gradient of the policy network at the previous tasks, $g_{\mathcal{M}}=\nabla_{s_{i}\sim\mathcal{M}}\mathcal{L}_{\pi_{\theta}}(s_{i})$.
@@
Algorithm 1 summarizes the generation of Adv-GEM examples. Adv-GEM starts with data collection by interacting with the environment and then adds the transitions $\{(s_{\mathcal{D}_{i}},a_{\mathcal{D}_{i}},r_{\mathcal{D}_{i}},s_{\mathcal{D}_{i}}^{\prime})\}$ to the episodic replay buffer $\mathcal{M}$. Via Adv-GEM augmentation, we obtain the augmentation samples according to Eq. 3. Finally, at the end of the task, we store $N_{\mathcal{M}}$ samples in the episodic memory $\mathcal{M}$. To verify the impact of the episodic memory size, we perform the ablation study at section 4.2. Via embedding the gradient of episodic memory into data augmentation and generating the adversarial examples, Adv-GEM can mitigate the inherent overfitting problem of the memory-based methods [20, 2].
@@ 3.3 Data Augmentation Framework of Continual RL
In continual RL, since an agent needs to interact with the different environments to collect new data samples, data augmentation serves as a simple technique that allows the agent to increase the diversity of data for each task. Data augmentation from states assume that input states expand to physically realizable in each task. Since the state and action spaces for each task in the continual RL remain the same, via data augmentation, we can increase the probability that samples from future tasks will be closer to the current data distribution.
@@ 4.2 Results
Effect of Data Augmentation. To validate the effect of data augmentation at MW4, we compare a base EWC agent and a base Packnet agent with various data augmentation methods. The average performance, forgetting, and forward transfer are tabulated in Table. 1. We found that RAS, state-switch, mixup, Adv-AUG, and Adv-GEM outperform the base methods on average performance and forward transfer both at EWC and PackNet. For the average performance, the EWC + Adv-GEM and PackNet + Adv-GEM methods are 0.74 and 0.88 respectively, suggesting their effectiveness over different continual RL methods. However, we also see that Uniform noise, Gaussian noise, and dimension-dropout can hurt the average performance of the continual RL agent when it is combined with EWC. For PackNet method, the Uniform noise and dimension-dropout augmentation decreased the average performance. These declined performances may be because they generate out-of-domain augmented data. The forward transfer represents the ability that continual methods to transfer knowledge forward. We found that RAS, mixup, Adv-AUG and Adv-GEM can highly increase the forward transfer both with EWC and PackNet. Moreover, the catastrophic forgetting of EWC + Adv-GEM and PackNet + Adv-GEM decrease to 0.06 and -0.10 respectively, which also validates that data augmentation can enhance the diversity of data distribution.
Figure 3 demonstrates that Adv-GEM augmentation significantly improves the average performance of both EWC and PackNet. EWC + Adv-GEM even outperforms PackNet, due to better test performance in tasks 1 and 3. PackNet + Adv-GEM surpasses PackNet with higher performance in task 4. These results clearly demonstrate the effectiveness of data augmentation in continual RL.
@@
Experiments at CW10. To assess the effectiveness of longer task scenarios, we evaluate the impact of data augmentations on CW10 environments. We apply all augmentation methods at the EWC and PackNet methods. As shown in Table 2, RAS, Adv-AUG, and Adv-GEM improve the average performance on CW10 when combined with EWC. Notably, the Adv-GEM transformation achieves a higher average performance value than the Adv-AUG transformation. In terms of catastrophic forgetting and forward transfer, both Adv-AUG and Adv-GEM have similar performance with the original EWC method. As for PackNet, both Adv-AUG and Adv-GEM outperformed the original method in terms of average performance. And the Adv-AUG + PackNet has 0.3 forward transfer value. These results showcase that data augmentations, like RAS, state-switch, mixup, Adv-AUG, Adv-GEM can improve the average performance of EWC and PackNet at MW4. And the result from CW10 validates the effectiveness of Adv-AUG and Adv-GEM for longer task scenarios.
Table 3: Average performance of different memory size for MW4.
@@
Epsilon. To investigate the effect of different epsilon values, which is the adversarial attack size, in Adv-AUG and Adv-GEM, we compare the value of 0.1, and 0.01 in MW4 respectively. As the scale of the input states, the epsilon value should be selected properly. Table. 4 shows that $\epsilon=0.1$ has better average performance. Due to the limitation of the computing source, we did not increase the epsilon to test. According to the scale of state input at the CW10 experiment, there is no need to increase the adversarial attack size over 1, which will be out of the physical limitation.
@@ 5 Conclusion
This paper proposes the data augmentation benchmark for the state-based continual RL. We find that data augmentations alone can significantly improve the average performance and forward transfer of continual RL methods operating from the state. And we introduce adversarial augmentation with memory gradient, which is to enhance the diversity of data distribution in continual RL by embedding the gradient episodic memory into adversarial examples. The proposed data augmentation framework is general, flexible, and easily expandable. For future work, we hope to increase the efficiency of adversarial augmentation and validate the data augmentation at more different continual RL settings and real-world applications.

### v3

@@ Abstract
Data efficiency of learning, which plays a key role in the Reinforcement Learning (RL) training process, becomes even more important in continual RL with sequential environments. In continual RL, the learner interacts with non-stationary, sequential tasks and is required to learn new tasks without forgetting previous knowledge. However, there is little work on implementing data augmentation for continual RL. In this paper, we investigate the efficacy of data augmentation for continual RL. Specifically, we provide benchmarking data augmentations for continual RL, by (1) summarising existing data augmentation methods and (2) including a new augmentation method for continual RL: Adversarial Augmentation with Gradient Episodic Memory (Adv-GEM). Extensive experiments show that data augmentations, such as random amplitude scaling, state-switch, mixup, adversarial augmentation, and Adv-GEM, can improve existing continual RL algorithms in terms of their average performance, catastrophic forgetting, and forward transfer, on robot control tasks. All data augmentation methods are implemented as plug-in modules for trivial integration into continual RL methods. Our code is available at https://github.com/WilliamWu96/DA_CRL_AGEM.
@@ 1 Introduction
  • We introduce a novel data augmentation method, Adv-GEM, specifically designed for continual RL. Our experimental results indicate that Adv-GEM achieves the highest positive impact on average performance compared to other augmentation techniques.
@@ 3 Data Augmentation for Continual RL
In this section, we introduce how to use data augmentation for continual RL. We start with the problem formulation in Section 3.1. Then, in Section 3.2, we introduce 8 different data augmentation methods. Finally, we present the general data augmentation framework for continual RL in Section 3.3.
@@ 3.2 Data Augmentation Methods
Figure 1: Framework for Adv-GEM data generation.
@@
Algorithm 1 summarizes the generation of Adv-GEM examples. Adv-GEM starts with data collection by interacting with the environment and then adds the transitions $\{(s_{\mathcal{D}_{i}},a_{\mathcal{D}_{i}},r_{\mathcal{D}_{i}},s_{\mathcal{D}_{i}}^{\prime})\}$ to the episodic replay buffer $\mathcal{M}$. Via Adv-GEM augmentation, we obtain the augmentation samples according to Eq. 3. Finally, at the end of the task, we store $N_{\mathcal{M}}$ samples in the episodic memory $\mathcal{M}$. To verify the impact of the episodic memory size, we perform the ablation study at section 4.2. Via embedding the gradient of episodic memory into data augmentation and generating the adversarial examples, Adv-GEM can mitigate the inherent overfitting problem of the memory-based methods [20, 2]. Adv-GEM offers several advantages in continual learning. By incorporating gradients from previous tasks, it encourages learning features useful across multiple tasks, leading to more generalized representations. This addresses the stability-plasticity dilemma by balancing adaptation to new information with retention of past knowledge. Adv-GEM’s adversarial examples expand the training distribution relevantly, guiding exploration towards robust solutions. By aligning generated examples with previous task gradients, it implicitly regularizes learning, reducing catastrophic forgetting. These insights suggest Adv-GEM produces robust, generalizable representations that maintain performance across tasks while avoiding overfitting, addressing a key challenge in continual RL.
@@ 3.3 Data Augmentation Framework of Continual RL
In continual RL, an agent must adapt to a sequence of tasks, each requiring interaction with a different environment to collect new data samples. Data augmentation serves as a powerful yet straightforward technique to enhance the diversity of data for each task. By applying augmentation to state representations, we expand the range of physically realizable inputs within each task’s domain. Given that the state and action spaces remain consistent across tasks in continual RL, data augmentation increases the likelihood that samples from future tasks will align more closely with the current task’s data distribution. This alignment facilitates better generalization and knowledge transfer between tasks, potentially mitigating catastrophic forgetting and improving overall performance in continual learning scenarios.
@@ 4.2 Results
Effect of Data Augmentation. To validate the effect of data augmentation at MW4, we compare a base EWC agent and a base Packnet agent with various data augmentation methods. The average performance, forgetting, and forward transfer are tabulated in Table. 1. We found that RAS, state-switch, mixup, Adv-AUG, and Adv-GEM outperform the base methods on average performance and forward transfer both at EWC and PackNet. For the average performance, the EWC + Adv-GEM and PackNet + Adv-GEM methods are 0.74 and 0.88 respectively, suggesting their effectiveness over different continual RL methods. However, we also see that Uniform noise, Gaussian noise, and dimension-dropout can hurt the average performance of the continual RL agent when it is combined with EWC. For PackNet method, the Uniform noise and dimension-dropout augmentation decreased the average performance. These declined performances may be because they generate out-of-domain augmented data. The forward transfer represents the ability that continual methods to transfer knowledge forward. We found that RAS, mixup, Adv-AUG and Adv-GEM can highly increase the forward transfer both with EWC and PackNet. EWC + Adv-GEM and PackNet + Adv-GEM reduce catastrophic forgetting to 0.06 and -0.10 respectively, demonstrating data augmentation’s ability to enhance data distribution diversity.
Figure 3 demonstrates that Adv-GEM augmentation significantly improves the average performance of both EWC and PackNet. EWC + Adv-GEM even outperforms PackNet, due to better test performance in tasks 1 and 3. PackNet + Adv-GEM surpasses PackNet with higher performance in task 4. Overall, these findings not only demonstrate the effectiveness of data augmentation in continual RL but also highlight the potential of Adv-GEM as a promising approach to addressing key challenges in the field, such as catastrophic forgetting and efficient knowledge transfer between tasks.
@@
Table 3: Average performance of different memory sizes for MW4.
@@
Experiments at CW10. To assess the effectiveness of longer task scenarios, we evaluate the impact of data augmentations on CW10 environments. We apply all augmentation methods at the EWC and PackNet methods. As shown in Table 2, RAS, Adv-AUG, and Adv-GEM improve the average performance on CW10 when combined with EWC. Notably, the Adv-GEM transformation achieves a higher average performance value than the Adv-AUG transformation. In terms of catastrophic forgetting and forward transfer, both Adv-AUG and Adv-GEM perform similarly to the original EWC method. As for PackNet, both Adv-AUG and Adv-GEM outperformed the original method in terms of average performance. The Adv-AUG + PackNet has 0.3 forward transfer value. These results showcase that data augmentations, like RAS, state-switch, mixup, Adv-AUG, Adv-GEM can improve the average performance of EWC and PackNet at MW4. The consistent improvement across EWC and PackNet algorithms, and MW4 and CW10 benchmarks, highlights the versatility and robustness of these data augmentation techniques. This suggests they address fundamental continual learning challenges like catastrophic forgetting and knowledge transfer, rather than exploiting specific algorithm or environment characteristics.
@@
Epsilon. To investigate the effect of different epsilon values, which is the adversarial attack size, in Adv-AUG and Adv-GEM, we compare the value of 0.1, and 0.01 in MW4 respectively. As the scale of the input states, the epsilon value should be selected properly. Table. 4 shows that $\epsilon=0.1$ has better average performance. Due to the limitation of the computing source, we did not increase the epsilon to test. According to the scale of state input at the CW10 experiment, there is no need to increase the adversarial attack size over 1, which would exceed the physical constraints of the system.
@@ 5 Conclusion
This paper introduces a data augmentation benchmark for state-based continual RL, showing significant improvements in performance and forward transfer. We present adversarial augmentation with memory gradient, enhancing data diversity through gradient episodic memory. Our flexible framework is easily expandable. Future work will focus on improving augmentation efficiency, validating the approach in diverse real-world applications, and developing adaptive strategies for varying task complexities.
Acknowledgements. We thank the anonymous reviewers for their thoughtful comments. XH’s contribution is supported by the UK EPSRC through End-to-End Conceptual Guarding of Neural Architectures [EP/T026995/1].